For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation?
4x − y = 8
2x + 3y = 9

Answer by ewatrrr(24785)

 
Hi
4x - y = 8 0r y+=+4x-8
2x + 3y = 9
2x + 3(4x-8) = 9
you substitute (4x-8) into the second equation?